You have been tasked with determining a fair stock price for Shire Brewing, a local craft beer manufacturer that will soon be going public for $22 per share. You are interested in investing in this company, but only if it is fairly priced (or undervalued). You have decided to use the free cash flow approach to determine the intrinsic price per share and have compiled the following data:

Free Cash Flow Data

Other Data

Year (t)

FCF

Growth rate of FCF beyond year 4 = 5%

1

$645,000

Weighted average cost of capital = 11%

2

775,000

Market Value of all debt = $2,300,000

3

930,000

Number of common shares outstanding = 500,000

4

1,115,000

Value of non-operating assets = $1,000,000

  1. What is the intrinsic price per share?
  2. Based on the calculated intrinsic price and the offer price of $22 should you buy the stock and why/why not?
  3. If the book value of equity is 10,000,000, what is the market value added of Company X?
Answer #1
FCF
1 645,000
2 775,000
3 930,000
4 1,115,000
TV 19,512,500
EV 15,478,070
Equity 14,178,070
Stock Price $      28.36

Terminal Value (TV) = FCF 4 x (1 + g) / (r - g) = 1,115,000 x (1 + 5%) / (11% - 5%) = $19,512,500

Enterprise Value (EV) = FCF1 / (1 + r) + FCF2 / (1 + r)^2 + FCF3 / (1 + r)^3 + (FCF4 + TV) / (1 + r)^4

= 645,000 / 1.11 + 775,000 / 1.11^2 + 930,000 / 1.11^3 + (1,115,000 + 19,512,500) / 1.11^4

= $15,478,070

Equity Value = EV - Debt + Non-operating assets

= 15,478,070 - 2,300,000 + 1,000,000

= $14,178,070

Stock Price = Equity Value / No. of shares = 14,178,070 / 500,000 = $28.36 ... a)

b) Yes. Because at $22 the stock is undervalued.